# SAR-Net: Multi-scale Direction-aware SAR Network via Global Information Fusion

### v2

@@ Abstract
Deep learning has driven significant progress in object detection using Synthetic Aperture Radar (SAR) imagery. Existing methods, while achieving promising results, often struggle to effectively integrate local and global information, particularly direction-aware features. This paper proposes SAR-Net, a novel framework specifically designed for global fusion of direction-aware information in SAR object detection. SAR-Net leverages two key innovations: the Unity Compensation Mechanism (UCM) and the Direction-aware Attention Module (DAM). UCM facilitates the establishment of complementary relationships among features across different scales, enabling efficient global information fusion. Among them, Multi-scale Alignment Module (MAM) and distinct Multi-level Fusion Module (MFM) enhance feature integration by capturing both texture detail and semantic information. Then, Multi-feature Embedding Module (MEM) feeds back global features into the primary branches, further improving information transmission. Additionally, DAM, through bidirectional attention polymerization, captures direction-aware information, effectively eliminating background interference. Extensive experiments demonstrate the effectiveness of SAR-Net, achieving state-of-the-art results on aircraft (SAR-AIRcraft-1.0) and ship datasets (SSDD, HRSID), confirming its generalization capability and robustness.
Index Terms:
Deep learning, SAR object detection, global fusion, direction-aware attention.
I Introduction
Object detection plays a pivotal role in remote sensing, with extensive applications in military operations, security surveillance, disaster assessment, and environmental monitoring[1, 2]. Synthetic Aperture Radar (SAR) object detection, a key technique in remote sensing, offers notable benefits in super-resolution imaging, hidden object detection, and all-weather surveillance, particularly in identifying aircraft and ships.
In recent years, with the rapid development of deep learning, an increasing number of researchers have turned their attention to design object detection methods using Convolutional Neural Networks (CNNs), yielding notable results [3, 4, 5, 6, 7, 8, 9]. However, the unique challenges posed by SAR images, characterized by their multi-scale and directional nature, coupled with abundant background details, complicate object detection tasks. These factors particularly hinder the detection of small objects amidst background noise, making the design of effective multi-scale object detection networks for SAR images a complex task. Consequently, developing strategies to accurately determine the direction and position of aircraft or ships in SAR images, despite background interference, stands as a critical challenge in this field.
@@ I Introduction
Current research in SAR image object detection primarily focuses on utilizing a FPN or its variants for feature fusion, offering a partial solution to multi-scale challenges. However, a significant limitation of these methods is their lack of global information integration at each layer. As illustrated in Fig. 1 (right side), local information flow between adjacent layers leads to a dilution of information across non-adjacent layers, resulting in information loss [14]. Recognizing the importance of global information, Bai et al. [15] proposed a Global Context-Guided Feature Balance Pyramid (GC-FBP), incorporating a global context module to bridge distant relationships between ship objects and their background. Yet, their approach of simply stacking context information onto the original feature map impedes effective interplay of information. Our method, depicted on the left side of Fig. 1, diverges from traditional algorithms that primarily depend on straightforward concatenation of adjacent feature layers. Instead, we adopt a global information fusion strategy, allowing for feature integration across various scales. This approach not only enhances the effectiveness and efficiency of the fusion process but also minimally impacts information loss during transmission.
@@
Traditional SAR object detectors, which predominantly use horizontal boxes, typically capture features of directional objects through an attention module [16] or by transforming features into the frequency domain [10]. However, these methods tend to neglect the correlation between the object and its contextual environment. This oversight often leads to ambiguous differentiation between the object and the background, with the resultant features being significantly noisy. Furthermore, these techniques generally overlook the crucial role of positional information. In contrast, our approach consider not only the object’s features but also the surrounding context. By integrating directional and positional information and embedding channel attention, our approach offers a comprehensive solution.
@@
We have distilled the problem of SAR image object detection into a broader challenge: addressing the issues of multi-scale and directional objects in predominantly background-dominated images. We base our considerations on three key premises: (1) the significant variation in object scales in SAR images. (2) the need to consider both shallow textural and deep semantic information of objects. (3) the directional nature of objects coupled with background interference. To address these challenges, we introduce a unity compensation mechanism, which includes three components: the Multi-scale Alignment Module (MAM), the Multi-level Fusion Module (MFM), and the Multi-feature Embedding Module (MEM). MAM initially unites multi-scale features across different spatial hierarchical levels. MFM then globally merges high-level semantic information with low-level detailed information, providing a comprehensive overview of the united features. Lastly, MEM compensates for any original feature to enrich their representational information. This multifaceted approach not only improves the detection of multi-scale objects but also significantly enhances information fusion capabilities.
@@
In addressing the third premise, we present a direction-aware attention module equipped with bidirectional attention polymerization capabilities. This innovative module employs deformable convolution and global pooling operations to extract directional and positional information along both vertical and horizontal axes of the input features. It then generates a direction-aware attention map, which is effectively integrated with the original input features. This approach adeptly tackles the challenge of accurately discerning directional and positional information amidst intricate background interference. Our contributions can be summarized in three primary points:
@@
$\bullet$ We introduce SAR-Net, a novel framework specifically designed for SAR image object detection. SAR-Net incorporates a unity compensation mechanism that globally unites multi-scale features of objects while compensating for the original feature. This ensures that global information is inherently integrated within features of different scales.
@@
The remainder of this paper is organized as follows. Section II reviews the development of object detection in the field of SAR images and related literature on solving multi-scale problems. Section III details our proposed framework and modules. In Section IV, the experimental setup is described, and qualitative and quantitative results are reported. Finally, conclusions are drawn in Section V.
II Related Work
II-A SAR Object Detection
Over the past few years, object detection in SAR images has aroused rising attention. Traditional detection algorithms, adhering to Constant False Alarm Rate (CFAR) principles [17], operate on a statistical model, dynamically setting the detection threshold based on a specified false alarm rate and the statistical distribution of background clutter. Consequently, a multitude of CFAR-based algorithms have surfaced. In a recent work [18], a novel decomposition approach was introduced to scrutinize the scattering between ships and the sea background. This decomposition technique, when integrated with CFAR, proves effective in ship detection. In a different vein, Lang et al. [19] enhanced the representation of pixels by incorporating spatial and intensity information, markedly improving the distinguishability between ships and background elements. However, the advent of high-resolution SAR images has introduced challenges such as complex ground interference and textured scenes accompanying multi-scale objects. This complexity renders the calculation of significant background clutter pixels in large-scale SAR images time-consuming. In response to the challenges of multi-scale object detection in SAR images, scholars have proposed several refined CFAR object detection algorithms. Examples include the scale sliding window object detection algorithm [20], the optimized super-pixel CFAR object detection algorithm [21], and the iterative CFAR object detection algorithm [22]. For detecting objects at larger scales, Zhai et al. [23] devised a ship detection algorithm emphasizing saliency and contextual information processing. This approach excels in focusing on sizable ships and background objects with prominent features, yielding commendable results. Despite these achievements, traditional methods still exhibit limitations in detecting small objects, necessitate manual parameter configuration, and heavily rely on manual design, presenting challenges for robustness and hindering their applicability to other SAR datasets. Consequently, there is a pressing need for more adaptive and automated approaches in SAR object detection.
@@ II-A SAR Object Detection
Owing to the inherent capacity for autonomous feature learning in deep learning, researchers have seamlessly integrated object detection in SAR images with deep learning methodologies. Shifting focus to SAR ship detection and identification, Zhang et al. [24] proposed a joint unsupervised super-resolution and SAR object detection network. This network augments object details while suppressing noise, providing valuable guidance to the detection network. In the realm of SAR aircraft detection and recognition, Zhao et al. [25] introduced an attention module, strategically fusing refined low-level texture features with high-level semantic features. This fusion strategy enhances aircraft detection rates by leveraging complementary information. Furthermore, Zhao et al. [26] devised a multi-branch dilated convolution feature pyramid method. This innovative approach minimizes redundancy while accentuating critical aircraft features through the establishment of dense connections. In a related endeavor, Zhao et al. [27] put forth a multitask learning framework for object detection (MLDet) tailored for ship detection in SAR images. This end-to-end framework encompasses tasks such as object detection, speckle suppression, and object segmentation. The collaborative execution of these tasks contributes to an enhanced detection performance. Addressing the specific challenge of large-scale SAR images, Yang et al. [28] concentrated on a multi-channel aggregation framework. This framework holistically considers both pre-processing and post-processing in algorithm optimization, leading to an overall improvement in the effectiveness of object detection in large-scale SAR images. The comprehensive strategy employed by these studies underscores the effectiveness of deep learning paradigms in advancing the state-of-the-art in SAR object detection and recognition.
@@
II-B Multi-scale Object Detection
Within the domain of deep learning, the application of image pyramids initially emerged as a solution to address multi-scale challenges in object detection, driven by its intuitive appeal [29, 30, 31, 32, 33]. Approaches such as SNIP [34] and SNIPER [35] stand out for their selection of specific scales for each resolution during multi-scale training. However, it’s noteworthy that the utilization of image pyramids in these approaches substantially escalates inference time, presenting a trade-off between accuracy and efficiency. An alternative strategy seeks to incorporate in-network feature pyramids, offering an approximation of image pyramids with reduced computational costs. Notably, the Single-Shot Detector (SSD) [36] exemplifies this concept by leveraging multi-scale feature maps from different layers. This enables the detection of objects spanning various scales at each feature layer. Despite the computational advantages, it is essential to acknowledge that bottom-up sampling methods, like those employed in SSD, may exhibit diminished accuracy when confronted with small objects. This limitation arises from the inherent inadequacy of semantic information in low-level features, particularly impacting the precision of small object detection. Balancing the trade-offs between computational efficiency and accuracy remains a critical consideration in the ongoing evolution of multi-scale object detection methodologies within the realm of deep learning.
To address the semantic limitations in low-level features, the Feature Pyramid Network (FPN) [11] was introduced. FPN employs a bottom-up pathway, a top-down pathway, and lateral connections to effectively merge features of diverse resolutions and scales. Building upon FPN, the Path Aggregation Network (PANet) [37] introduced a bottom-up path to enhance information fusion between different levels. Researchers at FSAF [38] identified a flaw in FPN’s heuristic-guided feature selection, highlighting the need for improvement. In contrast, FSAF extracts the feature representation of the region of interest (RoI) at a single feature level, overlooking any valuable information present on other feature levels of the pyramid. The SA-FPN [39], amalgamating Top-Down style FPN and Bottom-Up style FPN, combines the strengths of both, emerging as a more accurate module for perceiving scale variations. While previous methods integrated information at all levels through common operations like element-wise max or sum, these fusion mechanisms did not consider semantic differences among feature layers. EfficientDet [40] addressed this limitation by introducing a repeatable module (BiFPN) to enhance information fusion efficiency between different levels. YOLO-ASFF [41] employed an adaptive spatial feature fusion technique, learning weights during training to filter out inconsistencies. Recognizing that the feature maps in a pyramid predominantly consist of single-level features, M2Det [42] provided insights into this structure. Moreover, [43] extended FPN with the Asymptotic Feature Pyramid Network (AFPN) to facilitate interaction across non-adjacent layers. To overcome FPN’s limitations in detecting large objects, [44] proposed a parallel FPN structure for object detection with bi-directional fusion, while [45] introduced a refined FPN structure. SAFNet [46] introduced Adaptive Feature Fusion and Self-Enhanced Modules.
III Methods
This paper introduces SAR-Net, a SAR object detection framework based on a unity compensation mechanism. Built upon the YOLOv6 [47] architecture, SAR-Net consists of three components: backbone, neck, and head. Leveraging the advanced detection head of YOLOv6, we retain its head in the SAR-Net framework while focusing on innovations in the backbone and neck. In this section, we first present the overall architecture of SAR-Net, followed by an introduction to our proposed unity compensation mechanism, which addresses the issue of hindered information interaction due to local information fusion. Finally, we clarify the newly designed direction-aware attention module embedded in the backbone and neck, dedicated to acquiring object directional and positional information.
@@ III Methods
III-A Overall Framework of SAR-Net
The overall framework of SAR-Net is illustrated in Fig. 2, where the backbone and neck components are specifically tailored for SAR images. Specifically, considering that SAR image content is predominantly background, and detected objects are directional, using feature extraction networks designed for optical images directly cannot capture the directional and positional information of objects under complex background interference. Therefore, we have designed a direction-aware feature extraction backbone network, as shown in Fig. 3. It embeds our designed direction-aware attention module, reducing background interference while capturing fine-grained information, and providing effective assistance for the subsequent global fusion of multi-scale features in SAR images. Additionally, to better address the multi-scale challenge in SAR images and avoid information dilution caused by traditional FPN local fusion, we propose a unity compensation mechanism. This mechanism comprehensively fuses deep and shallow features at different scales, enabling each layer to focus on global information and providing robust prior knowledge for the detection head.
III-B Unity Compensation Mechanism
In our implementation, the process of unity and compensation corresponds to three modules: the Multi-scale Alignment module, the Multi-level Fusion module, and the Multi-feature Embedding module. These modules collaborate to unitedly compensate for features, we illustrate the detailed information in the following subsections. As shown in Fig. 2, the input to the neck includes the feature maps $C2$, $C3$, $C4$, and $C5$ extracted from the backbone, where $Ci\in\mathbb{R}^{N\times\tilde{C_{Ci}}\times H_{Ci}\times W_{Ci}}$. The batch size is denoted by $N$, the channels by $C$, and the sizes by $H\times W$. Additionally, the feature map sizes decrease progressively from $C2$ to $C5$, while the channel numbers double at each level.
III-B1 Multi-scale Alignment Module
Given the crucial premise that both shallow texture and deep semantic information of objects in SAR images are important, we choose the outputs $C2$, $C3$, $C4$, and $C5$ from the backbone for global feature alignment. We take two factors into consideration for aligning feature size: the computational complexity of global fusion and the necessity of preserving the fine-grained information extracted by the backbone. The former tends to favor smaller feature maps to reduce computational load, while the latter prefers larger feature maps to preserve detailed information. To balance these two aspects, the second smallest feature size is chosen as aligned size when four features need to be aligned while the smallest feature size is selected when aligning three features.

### v4

# Multi-scale direction-aware SAR object detection network via global information fusion

@@ Abstract
Deep learning has driven significant progress in object detection using Synthetic Aperture Radar (SAR) imagery. Existing methods, while achieving promising results, often struggle to effectively integrate local and global information, particularly direction-aware features. This paper proposes SAR-Net, a novel framework specifically designed for global fusion of direction-aware information in SAR object detection. SAR-Net leverages two key innovations: the Unity Compensation Mechanism (UCM) and the Direction-aware Attention Module (DAM). UCM facilitates the establishment of complementary relationships among features across different scales, enabling efficient global information fusion and transmission. Additionally, DAM, through bidirectional attention polymerization, captures direction-aware information, effectively eliminating background interference. Extensive experiments demonstrate the effectiveness of SAR-Net, achieving state-of-the-art results on aircraft (SAR-AIRcraft-1.0) and ship datasets (SSDD, HRSID), confirming its generalization capability and robustness.
keywords:
Deep learning , SAR object detection , Global fusion , Direction-aware attention
[…]
{graphicalabstract}
{highlights}
A SAR object detector to tackle information fusion and direction-aware problems.
Unity Compensation Mechanism (UCM) and Direction-aware Attention Module (DAM).
UCM unites multi-scale features and compensates for the original features.
DAM obtains positional and directional information in complex background.
We achieve SOTA accuracy among SAR-AIRcraft-1.0, SSDD and HRSID datasets.
1 Introduction
Object detection plays a pivotal role in remote sensing, with extensive applications in military operations, security surveillance, disaster assessment, and environmental monitoring [1]. Synthetic Aperture Radar (SAR) object detection, a key technique in remote sensing, offers notable benefits in super-resolution imaging, hidden object detection, and all-weather surveillance, particularly in identifying aircraft and ships.
In recent years, with the rapid development of deep learning, an increasing number of researchers have turned their attention to design object detection methods using Convolutional Neural Networks (CNNs), yielding notable results. However, the unique challenges posed by SAR images, characterized by their multi-scale and directional nature, coupled with abundant background details, complicate object detection tasks. These factors particularly hinder the detection of small objects amidst background noise, making the design of effective multi-scale object detection networks for SAR images a complex task. Consequently, developing strategies to accurately determine the direction and position of aircraft or ships in SAR images, despite background interference, stands as a critical challenge in this field.
@@ 1 Introduction
Current research in SAR image object detection primarily focuses on utilizing FPN or its variants for feature fusion, offering a partial solution to multi-scale challenges [4, 5]. However, a significant limitation of these methods is their lack of global information integration at each layer. As illustrated in Fig. 1 (right side), local information flow between adjacent layers leads to a dilution of information across non-adjacent layers, resulting in information loss [6]. Recognizing the importance of global information, Bai et al. [7] proposed a Global Context-Guided Feature Balance Pyramid (GC-FBP), incorporating a global context module to bridge distant relationships between ship objects and their background. Yet, information flows similar to FPN are still used, and their approach of simply stacking context information onto the original feature map impedes effective interplay of information. Our method, depicted on the left side of Fig. 1, diverges from traditional algorithms that primarily depend on straightforward concatenation of adjacent feature layers. Instead, we adopt a global information fusion strategy, allowing for feature integration across various scales. This approach not only enhances the effectiveness and efficiency of the fusion process but also minimally impacts information loss during transmission.
@@
Traditional SAR object detectors, which predominantly use horizontal boxes, typically capture features of directional objects through an attention module [8] or by transforming features into the frequency domain [2]. However, these methods tend to neglect the correlation between the object and its contextual environment. This oversight often leads to ambiguous differentiation between the object and the background, with the resultant features being significantly noisy. Furthermore, these techniques generally overlook the crucial role of positional information. In contrast, our approach considers not only the object’s features but also the surrounding context. By integrating directional and positional information and embedding channel attention, our approach offers a comprehensive solution.
@@
We have distilled the problem of SAR image object detection into a broader challenge: addressing the issues of multi-scale and directional objects in predominantly background-dominated images. We base our considerations on three key premises: (1) the significant variation in object scales in SAR images. (2) the need to consider both shallow textural and deep semantic information of objects. (3) the directional nature of objects coupled with background interference. To address these challenges, we introduce a unity compensation mechanism, which includes three components: the Multi-scale Alignment Module (MAM), the Multi-level Fusion Module (MFM), and the Multi-feature Embedding Module (MEM). MAM initially unites multi-scale features across different spatial hierarchical levels. MFM then integrates overarching semantic cues before globally amalgamating them with intricate, fine-grained details, providing a comprehensive overview of the united features. Lastly, MEM compensates for any original feature to enrich their representational information. This multifaceted approach not only improves the detection of multi-scale objects but also significantly enhances information fusion capabilities.
@@
In addressing the third premise, we present a direction-aware attention module equipped with bidirectional attention polymerization capabilities. This innovative module employs deformable convolution and global pooling operations to extract object positional and directional information along the rows and columns of input features. It then generates a direction-aware attention map, which is effectively integrated with the original input features. This approach adeptly tackles the challenge of accurately discerning directional and positional information amidst intricate background interference. Our contributions can be summarized in three primary points:
@@
$\bullet$ We present SAR-Net, an innovative framework tailored specifically for object detection in SAR images. SAR-Net incorporates a unity compensation mechanism that globally unites multi-scale features of objects while compensating for the original feature. This ensures that global information is inherently integrated within features of different scales.
@@
The subsequent sections of this paper are structured as follows: Section II provides an overview of the evolution of object detection within the realm of SAR imagery, alongside a survey of relevant literature addressing multi-scale challenges. Section III details our proposed framework and modules. In Section IV, the experimental setup is described, and qualitative and quantitative results are reported. Finally, conclusions are drawn in Section V.
2 Related work
2.1 SAR object detection
Object detection in SAR images has gained significant attention in recent years. Traditional detection algorithms, which follow the Constant False Alarm Rate (CFAR) principles [9, 10, 11], operated based on a statistical model. In prior works [12], researchers aimed to segregate objects from the background to emphasize disparities and streamline detection processes. Some other works [13] were devoted to solving the problem of multi-scale object detection in high-resolution images. Despite these achievements, traditional methods still face constraints in detecting small objects and heavily depend on manual design, posing challenges for robustness and limiting their applicability to different SAR datasets. Hence, there is a critical demand for more adaptable and automated approaches in SAR object detection.
@@ 2.1 SAR object detection
Owing to the inherent capacity for autonomous feature learning in deep learning, researchers have seamlessly integrated object detection in SAR images with deep learning methodologies [14]. Shifting focus to SAR ship detection and identification, Zhang et al. [15] proposes a joint unsupervised super-resolution and SAR object detection network. This network augments object details while suppressing noise, providing valuable guidance to the detection network. In the realm of ship detection and recognition, Guo et al. [14] developed three modules for feature refinement and fusion, while utilizing a head module to reduce the impact of background noise. Furthermore, Li et al. [2] utilized the FPN to create spatial multi-scale features, specifically for ship objects, through a top-down hierarchical approach. In a related endeavor, Li et al. [5] introduced the AFF module to strategically integrate local shallow features into adjacent layers. This self-decision approach enables the flexible integration of different features, enhancing the model’s ability to extract object features while also providing a certain level of robustness. Addressing the specific challenge of large-scale SAR images, Yang et al. [16] holistically considers both pre-processing and post-processing in algorithm optimization, leading to an overall improvement in the effectiveness of object detection in large-scale SAR images. The comprehensive strategy employed by these studies underscores the effectiveness of deep learning paradigms in advancing the state-of-the-art in SAR object detection and recognition.
@@
2.2 Multi-scale object detection
In the realm of deep learning, the adoption of image pyramids initially emerged as a solution for tackling multi-scale challenges in object detection, driven by its intuitive appeal. Notable approaches distinguished themselves by selecting specific scales for each resolution during multi-scale training. However, it’s essential to note that these methods significantly increase inference time, creating a trade-off between accuracy and efficiency. Another approach is to build a feature pyramid inside the network, providing an approximate image pyramid and reducing computational costs. This enables the detection of multi-scale objects at each feature layer but can lead to decreased accuracy in detecting small objects. The issue arises from the low-level features themselves lacking semantic information, which is particularly crucial for small object detection.
To address the semantic limitations in low-level features, the Feature Pyramid Network (FPN) [3] is introduced. FPN offers an efficient structural design for merging multi-scale features via skip connections and local information transfer. Researchers at FSAF [17] identify a flaw in FPN’s heuristic-guided feature selection, highlighting the need for improvement. In contrast, FSAF focuses attention on regions of interest within each hierarchical feature level, aiming to extract features from a single level, but ignores valuable information between different hierarchical levels. The SA-FPN [18], amalgamating Top-Down style FPN and Bottom-Up style FPN, combined the strengths of both, emerging as a more accurate module for perceiving scale variations. While previous methods integrate information at all levels through common operations like element-wise max or sum, these fusion mechanisms do not consider semantic differences among feature layers. EfficientDet [19] addressed this limitation by introducing a repeatable module (BiFPN) to enhance information fusion efficiency between different levels. Recognizing that the feature maps in a pyramid predominantly consist of single-level features, M2Det [20] provided insights into this structure. Moreover, [21] extended FPN with the Asymptotic Feature Pyramid Network (AFPN) to facilitate interaction across non-adjacent layers. To overcome FPN’s limitations in detecting large objects, [22] proposed a parallel FPN structure for object detection with bi-directional fusion.
@@ 2.2 Multi-scale object detection
3 Methods
This paper presents SAR-Net, an SAR object detection framework that incorporates a unity compensation mechanism and a direction-aware attention module. Built upon the YOLOv6 [23] architecture, SAR-Net consists of three components: backbone, neck, and head. Leveraging the advanced detection head of YOLOv6, we retain its head in the SAR-Net framework while focusing on innovations in the backbone and neck. Within this section, we first present the comprehensive architecture of SAR-Net, followed by an introduction to our proposed unity compensation mechanism, which addresses the issue of hindered information interaction due to local information fusion. Finally, we clarify the newly designed direction-aware attention module embedded in the backbone and neck, dedicated to acquiring object directional and positional information.
@@ 3 Methods
3.1 Overall framework of SAR-Net
The comprehensive framework of SAR-Net is depicted in Fig. 2, where the backbone and neck components are specifically tailored for SAR images. Specifically, considering that SAR image content is predominantly background, and detected objects are directional, using feature extraction networks designed for optical images directly cannot capture the directional and positional information of objects under complex background interference. To this end, we have designed a direction-aware feature extraction backbone network, as shown in Fig. 3. It embeds our designed direction-aware attention module, reducing background interference while capturing fine-grained information, and providing effective assistance for the subsequent global fusion of multi-scale features in SAR images. Additionally, to better tackle the challenge of multi-scale in SAR images and avoid information dilution caused by traditional FPN local fusion, we propose a unity compensation mechanism. This mechanism comprehensively fuses low-level and high-level features across various scales, enabling each layer to focus on global information and providing robust prior knowledge for the detection head.
3.2 Unity compensation mechanism
In our implementation, the unity and compensation process is facilitated by three modules: the Multi-scale Alignment module, the Multi-level Fusion module, and the Multi-feature Embedding module. These modules collaborate to unitedly compensate for features, we illustrate the detailed information in the following subsections. As depicted in Fig. 2, the input to the neck includes the feature maps $C2$, $C3$, $C4$, and $C5$ extracted from the backbone, where $Ci\in\mathbb{R}^{N\times\tilde{C_{Ci}}\times H_{Ci}\times W_{Ci}}$. The batch size is represented as $N$, the number of channels as $C$, and the dimensions as $H\times W$. Additionally, the feature map sizes decrease progressively from $C2$ to $C5$, while the channel numbers double at each level.
3.2.1 Multi-scale alignment module
Given the crucial premise that both shallow texture and deep semantic information of objects in SAR images are important, we choose the outputs $C2$, $C3$, $C4$, and $C5$ from the backbone for global feature alignment. We consider two factors for aligning feature size: the computational complexity of global fusion and the necessity of preserving the fine-grained information extracted by the backbone. The former tends to favor smaller feature maps to reduce computational load, while the latter prefers larger feature maps to preserve detailed information. To balance these two aspects, the second smallest feature size is chosen as aligned size when four features need to be aligned while the smallest feature size is selected when aligning three features.